function are                                                                                                                        

P1 = 80 – 2.5Q1

P2 = 180 – 10Q2

The monopolist total cost

TC = 50 + 40Q

Determine the price to be charged in the two markets and amount of output to be sold in each market so that profit is maximize. find the total profit to be made from the strategy of price discrimination.


1
Expert's answer
2021-04-29T10:30:16-0400
"Q_1=320-4p_1"

"Q_2=18-0.1p_2"

"Q=Q_1+Q_2"

"Q=338-4p_1-0.1p_2"

"MC_1=MC_2=40"

"TR_1=p_1Q_1=320p_1-4p_1^2"

"MR_1=320-8p_1"

"p_1=35"

"Q_1=180"

"TR_2=p_2Q_2=18p_2-0.1p_2^2"

"MR_2=18-0.2p_2"

"p_2=1.1"

"Q_2=16.9"